OpenStudy (anonymous):

Factorise: x^2 + 3x - 5x - 15 ?

OpenStudy (anonymous):

(x+ 3) (x-5)

OpenStudy (anonymous):

\[x ^{2}+3x-5x-15=x(x+3)-5(x+3)=(x+3)(x-5)\]